If an investor purchases a stock that appreciates from $50 per share to $100 per share over that twelve-month period, what is his/her return with 50% margin (assume no interest rate)? What is the return assuming no margin?
Answer #1

HI,

A return on stock will be selling price upon investment.

Here initial price of stock = $50

margin =50%

so investor initial investment = 50*50% = $25

Selling price = $100

so Profit = $100-50 = $50

Return on investment% = Profit*100/initial investment = 50*100/25 = 200%

Assuming no margin initial investment= $50

So Return on investment% = 50*100/50 = 100%

Thanks